OpenStudy (lolacole12):

3+(7+4)x 9 ---------- 2+7x3 @gabbyalicorn

OpenStudy (gabbyalicorn):

3+(7+4)(x)(9) =3+99x Answer: =99x+3 2+7x^3 Answer: =7x^3+2 =37x3+99x+2

OpenStudy (lolacole12):

oh